Fill in the correct number to complete correctly the following statements:

Trifolium hybridum is a diploid species with 16 chromosomes per somatic cell.

            a) A tetrasomic of this species would have a somatic chromosomes number of ______.

            b) A triploid would have _____ chromosomes.

            c) A trosomic would have _____ chromosomes.

            d) An autotetraploid would have _____ chromosomes.

            e) A monosomic would have _____ chromosomes.
Answer #1

Answer:

a) A tetrasomic of this species would have a somatic chromosomes number of __18____.

Explanation:

Tetrasomy = 2n+2

= 16+2

= 18

b) A triploid would have _24____ chromosomes.

Explanation:

2n = 16

n = 16/2=8

Triploid = 3n = 3 * 8 = 24

c) A trosomic would have _17____ chromosomes.

Explanation:

Trisomy = 2n+1

= 16+1

= 17

d) An autotetraploid would have _32____ chromosomes.

Explanation:

2n = 16

n = 16/2=8

Autotetraploid = 4n = 4 * 8 = 32

e) A monosomic would have __15___ chromosomes.

Explanation:

Monosomy = 2n-1

= 16-1

= 15